# Containerization in Multi-Cloud Environment: Roles, Strategies, Challenges, and Solutions for Effective Implementation

### v1

@@ 3.6. Automation Challenges and Solution Framework (RQ4 and RQ5)
S27 Multi-Cloud Deployment and Monitoring Tackling challenges in automating deployment and monitoring of applications across multi-cloud platforms Multi-cloud Strategy Multi-cloud strategy for risk distribution and provider dependency reduction in multi-cloud environment
S06 Container Placement Automation Addressing the complexity of automating container placement in multi-cloud environment Intelligent Container Placement Develop a machine learning-based optimization algorithm for intelligent container placement across multiple cloud environments, focusing on load balancing and minimizing network latency.
S07 Application Runtime Switch Automation Enabling automated application runtime switch among various cloud services Service Driver Service driver, multi-cloud application service access model, runtime service update, development framework in multi-cloud environment
S15 Fine-Grained Automation at Component Level Implementing autonomic fine-grained solutions at the service component level for better automation Mobility Service and Replication Mobility service, migration and replication of critical software components, autonomic fine-grained service migration/replication in multi-cloud environment
S28 DevOps Automation Ensuring effective automation of software development and operations management DevOps Development Model DevOps development model, continuous development, automated tools and management systems, autonomic management system, performance models, systematic decision-making in multi-cloud environment
S29 Service Deployment Automation Addressing manual deployment issues by automating service deployment Automation Scripts Automation scripts, automated deployment, and error reduction in multi-cloud environment
S30 Deployment Time Security Automation Implementing security automation during deployment time Automated Security Measures Automated security measures, security evaluation and bootstrap, vulnerability scanning, firewall deployment, Chef integration in multi-cloud environment
S31 Distributed Simulation Automation Automating distributed simulation environments in cloud infrastructures DevOps Methodologies DevOps methodologies, automated developer operations, rapid microservices deployment, containerization paradigm, DEVS Simulation Farm, Docker technology, SOA Microservices paradigm, DEVSML Stack in multi-cloud environment
S32 Software Deployment Speed Improvement Addressing slow software deployment issues by implementing effective automation, enhancing Quality of Experience (QoE) Reorganized Container Images Reorganized container images, modified Docker deployment process, asynchronous image layer retrieval, profiled container execution, overlay file system in multi-cloud environment
S36 Container Startup Automation Identifying essential files for container startup automation Profiled Container Execution Profiled container execution, essential file determination, special image layer creation in multi-cloud environment
S38 Dynamic Scaling of Resources Automating frequent deployment and dynamic scaling of resources in multi-cloud environment for improved efficiency and performance CSP Automation Capability CSP automation capability for user agility in containerization of multi-cloud environment
S24 Container Deployment and Scaling Implementing automated container deployment and scaling based on resource demand across multiple cloud platforms Kubernetes Kubernetes for automated container deployment and multi-cloud scaling in multi-cloud environment
S48 Elastic Resource Provisioning Implementing elastic resource provisioning for fluctuating business workloads in a containerized multi-cloud environment through automation Elastic Resource Provisioning Design a predictive analytics system coupled with automation tools for real-time, elastic resource provisioning, addressing fluctuating business workloads in a containerized multi-cloud environment
S50 VF Placement and Service Deployment Automating optimal Virtual Function (VF) placement and dynamic service deployment in multi-cloud environment VF Placement Optimization Optimization model for virtual function (VF) placement and dynamic VF allocation in multi-cloud environment
S68 Container and VM Deployment Orchestrating and automating the deployment of containers and VMs in multi-cloud environment Multi-Cloud Orchestration Develop a comprehensive multi-cloud management platform for streamlining orchestration and automation of containers and virtual machine deployment across different cloud environments
@@
S04 Resource Overbooking Management Automating the management of resource overbooking in service containers across multi-cloud platforms ML-based Resource Overbooking Detection Machine learning-based resource overbooking detection in multi-cloud environment
S21 VCluster Management Automating cluster creation, deployment, and management for multiple Virtual Clusters (Vclusters) across diverse application frameworks ClaaS Service Model ClaaS service model, virtualized cluster environment, easy deployment and management in multi-cloud environment
S39 Containerized Applications Management Automating deployment, scheduling, and management of containerized applications in multi-user and multi-cloud environment Custom Kubernetes Scheduler Custom Kubernetes scheduler (label-affinity-scheduler) for container orchestration in multi-cloud environment
S41 VM Types Selection in Data Centers Addressing the large and complex search space of different VM types in various data centers through automation GA-based Algorithm GA-based algorithm, domain-tailored service clustering method for complexity reduction in multi-cloud environment
S42 Resource Commissioning and De-Commissioning Speeding up resource commissioning and de-commissioning through automation for timely system scalability MicroCloud Architecture MicroCloud architecture, fine-grained resource allocation, TOSCA library, meta-workflow, coordinated adaptation workflow in multi-cloud environment
S43 Cloud Federation Automation Implementing automation support for federation or inter-platform portability in multi-cloud environment Middleware Platform Middleware platform for security, cost, and performance optimization, data locality, privacy needs in multi-cloud and MEC environments
S45 Resource Provisioning and Elasticity Detection Automating resource provisioning for cloud services, monitoring runtime configurations, and detecting elasticity actions JCatascopia JCatascopia, a fully-automated and modular cloud monitoring framework with elasticity actions, metric rule mechanism, and low-level monitoring metrics in multi-cloud environment
S51 Resource Acquisition Automation Automating resource acquisition processes for applications to function without human intervention Aneka Platform Aneka platform for automatic acquisition and release of resources from multiple clouds in multi-cloud environment
S74 SLA Enforcement Automation Automating the testing and corrective actions for SLA enforcement in containerized multi-cloud environment to reduce expense and time consumption SLA Compliance Testing Manual testing and corrective actions for SLA compliance are expensive and time-consuming in multi-cloud environment
@@
S01 Data Integrity during Migration Safeguarding data integrity during the automation of migration processes in multi-cloud environment Autonomous Management Cycle Autonomous management cycle for efficient deployment in multi-cloud environment
S17 Legacy Web Applications Migration Facilitating the automated migration of legacy web applications to cloud services Reusable Architectural Pattern Reusable architectural pattern for containerization in multi-cloud environment
S69 Application Porting to Cloud Automating the porting of existing applications to cloud infrastructures CloudSME Simulation Platform CloudSME Simulation Platform for fast development and deployment of complex multi-cloud applications, commercial accounting, and billing
Category 5: Testing and Banchmakring Automation
S09 User-Oriented Testing Automation Accelerating user-oriented software testing in diverse and heterogeneous environments through automation S Container-based Lightweight Virtualization Container-based lightweight virtualization, just-in-time deployment, streamlined multicomponent prototyping in multi-cloud environment
S02 Benchmarking Automation Implementing an automated benchmarking process for cloud hosts from diverse providers in multi-cloud environment Smart Docker Benchmarking Orchestrator Smart Docker Benchmarking Orchestrator for containerization in multi-cloud environment
S13 CI for Dependable Automation Leveraging Computational Intelligence (CI) to enhance dependability in automated multi-cloud containerized environments Dependability Enhancement with CI Leverage Computational Intelligence (CI) to enhance dependability in automated multi-cloud containerized environments
@@
S05 Standardization in Automation Implementing standardization in the automation processes for containerized applications across multiple cloud environments Service Mesh and Multi-cloud Federation Service mesh network and multi-cloud cluster federation for containerization in multi-cloud environment
S70 Cloud Provider Selection Developing optimal strategies for cloud provider and node configuration selection to balance cost and runtime in a multi-cloud, containerized environment Optimal Cloud Provider Selection Create an AI-driven decision-support tool that automatically selects the optimal cloud provider and node configuration, considering both cost-effectiveness and runtime efficiency in a multi-cloud, containerized environment.
S16 Unified Template for Microservices Addressing the lack of a unified template and specification for automating microservices deployment across multiple cloud platforms Tool Service Model Tool service model, invocation relationships, service function chains, objective function optimization, accelerated Particle Swarm Optimization, configurable objective parameters in multi-cloud environment
S64 Cloud Interoperability Addressing the interoperability issues between different clouds through automation multi-cloud Harmony Develop an AI-driven multi-cloud management platform to automate cross-cloud interoperability, ensuring seamless data transfer and communication between disparate cloud environments
S63 Cloud Applications Standardization Promoting the adoption of worldwide standards or interfaces for cloud applications and their management through automation Cloud Service Interoperability Cloud service providers and middleware for interoperability, standardized interfaces, and hybrid solutions in multi-cloud environment
S18 Package Creation and Scaling Automating the creation and scaling of packages using VM and Docker based on machine-readable definition files Package Automation with VM Docker Implement a scriptable automation framework that facilitates creation, scaling, and management of packages using VM and Docker based on machine-readable definition files
@@
S55 5G Ecosystem Component Selection Simplifying the complexity of 5G ecosystem component selection through automation Open-source Software Components Open-source software components for customizable container and virtual machine orchestration in contemporary deployments
S62 End-to-End Tail Latency Management Automating end-to-end tail latency management in a microservice architecture Performance Modeling Approach Modeling approach for performance evaluation of container-level and VM-level data in multi-cloud environment
S79 Distributed Application Complexity Addressing the increasing complexity of managing distributed applications in containerized multi-cloud environment Complexity of Distributed Applications Increasing complexity of distributed applications and operating multi-cloud management topologies involving CMPs
@@
S72 Network Complexity Managing the complexity of network control and administration within containerized applications across multiple cloud environments. Network Management Complexity Complexity of network management and control in multi-cloud environment
S83 Social Network Service Discovery Improving service discovery algorithms for tagging in online social networks within containerized multi-cloud environment Automated Discovery and Selection Automating discovery and selection of relevant and appropriate services for composition and QoS management in multi-cloud environment
S08 Complex Application Administration Handling growing application complexity due to new technologies and needs through automation Autonomic Computing Systems Autonomic Computing Systems with knowledge transfer for containerization in multi-cloud environment
@@ 3.7. Deployment Challenges and Solution Framework (RQ4 and RQ5)
Table 13 presents a comprehensive overview of the challenges and corresponding solutions faced during the deployment of containerized applications in multi-cloud environment. As organizations increasingly adopt containerization and multi-cloud strategies to enhance scalability, flexibility, and resource utilization, they encounter various obstacles that require effective solutions. The figure consists of nine categories, each representing a specific deployment challenge, along with the corresponding challenge description and its solution. These categories encompass a wide range of issues, including deployment complexity and Orchestration, access and communication control, security and compliance management, multi-cloud deployment coordination and Integration, monitoring and scalability challenges, microservices architecture and containerization, and network connectivity and hybrid cloud integration challenges.
Deployment Complexity and Orchestration: This category focuses on addressing challenges and providing solutions related to the complexity and orchestration of container deployment in multi-cloud environment. One of the significant challenges faced in this context is Performance Testing Deployment. It involves evaluating testing procedures and assessing performance metrics during container deployment across multiple cloud environments. To overcome this challenge, the corresponding solution is Standardized Benchmarking and Automation. This solution entails adopting Standardized Benchmarking and automatic systems to ensure consistent and reliable performance evaluation. Another challenge is Deployment Validation, which requires conducting thorough validation and artifact checks during container deployment in multi-cloud environment. To effectively address this challenge, the suggested solution is the implementation of a Testing Process Management System. This system enhances and oversees the validation process for deployment, making sure that all essential criteria are met before deployment proceeds. The challenges outlined in this category can simplify and optimise container deployment in complex multi-cloud environment, facilitating efficient and reliable operations across different cloud service providers.
@@
Access and Communication Control: This category emphasizes maintaining secure multi-cloud connectivity and managing user access control. Challenges in this category include ensuring secure access to resources and secure communication between containers and across clouds. The proposed solutions involve implementing encryption and authentication mechanisms, using secure communication protocols, and executing operations in a provider-agnostic manner to maintain security.
@@
Security and Compliance Management in Multi-cloud Deployment: This category focuses on security and compliance challenges faced during multi-cloud container deployment. One challenge, Vendor Lock-in Prevention Deployment, aims to prevent vendor lock-in and ensure modular and loosely-coupled deployment in multi-cloud environment. The solution proposed is the Cloud Modelling Framework, which introduces a framework for specifying provisioning and deployment to enhance compatibility with existing ACSs and cloud solutions. Another challenge in this category is Security Deployment in Multi-container Environment, which addresses security and privacy concerns during multi-container deployment on the same OS kernel in multi-cloud environment. The solution is Cross-Container Isolation, which enforces cross-container isolation to enhance security.
Multi-cloud Deployment Coordination and Integration: This category deals with challenges related to coordination and integration during multi-cloud deployment. One challenge, Multi-cloud PaaS Deployment, aims to overcome high entry barriers for deploying a multi-cloud PaaS infrastructure. The proposed solution is Lightweight Proxies for PaaS Integration, which utilizes lightweight proxies to facilitate individual PaaS services integration. Another challenge is IoT Application Deployment, which addresses challenges related to deploying IoT applications in multi-cloud environment. The solution is based on the DRA Framework and CI Broker for Multi-Cloud Deployment, which leverages the DRA framework and CI broker component for efficient multi-cloud deployment.
Monitoring and Scalability Challenges: This category focuses on challenges related to monitoring, performance, and scalability during multi-cloud container deployment. For instance, Legacy Code Migration Deployment deals with handling legacy code migration, tenant engine separation, and composition modeling during container deployment in multi-cloud environment. The proposed solution is the Reusable Architectural Pattern with Docker and Cloudant, which advocates for adopting a reusable architectural pattern with Docker containers and Cloudant persistence layer.
Microservices Architecture and Containerization: This category addresses challenges associated with microservices architecture and containerization in multi-cloud environment. For example, Autonomic System Deployment aims to implement an autonomic management system for deploying multi-layer, multi-tier web architectures in multi-cloud environment. The solution proposed is the Self-Tuning Performance Model and Autonomic Management, which introduces a self-tuning performance model and autonomic management system.
Network Connectivity and Hybrid Cloud Integration: This category focuses on network connectivity and hybrid cloud integration challenges during multi-cloud deployment. One challenge, Inter-edge Bandwidth Deployment, aims to address challenges related to interconnecting distributed localities and provisioning inter-edge bandwidth.
Cloud Deployment Constraints and Challenges: This category deals with various constraints and challenges encountered during cloud deployment. For instance, ACS Compatibility Deployment aims to ensure compatibility with existing ACSs and cloud solutions during container deployment in multi-cloud environment. The solution proposed is the Scalable Multi-Cloud Deployment Framework, which proposes a scalable deployment framework for multi-cloud applications.
Infrastructure Provisioning and Container Deployment Challenges: This category addresses challenges related to infrastructure provisioning and container deployment in multi-cloud environment. For example, Fog Computing Deployment deals with fog computing hardware and container deployment time challenges. The proposed solution is Reorganized Container Images and Docker Deployment, which involves reorganizing container images and modifying the Docker deployment process.
Application Deployment Workflow and Orchestration Challenges: This category focuses on challenges related to application deployment workflow and orchestration in multi-cloud environment. One challenge is API Management Deployment, which deals with effectively managing different cloud providers’ APIs/interfaces for multi-cloud container deployment. The solution proposed is DevOps Approach for Multi-Cloud Applications, which advocates for adopting a DevOps approach for managing the lifecycle of multi-cloud applications.
@@
Table 13. Deployment challenge and solution framework for containerized applications in multi-cloud environment
ID Challenge Challenge Description Solution Solution Description
@@
S01 Performance Testing Deployment Evaluating testing and performance during container deployment in multi-cloud environment Standardized Benchmarking and Automation Adoption of standardized benchmarking applications and automatic systems
S09 Deployment Validation Conducting rigorous validation and artifact checks during container deployment in multi-cloud environment Testing Process Management System Implementation of a testing process management system
S02 Cost-effective Deployment Ensuring cost-effective and flexible container deployment options in multi-cloud environment Containerization and Microservice Architecture Implementation of containerization and microservice architecture
S38 Resource Management Deployment Managing diverse resources across multiple cloud service providers during container deployment User Preference-Based Multi-Cloud Resource Brokering Adoption of multi-cloud resource brokering dependent on user preferences and application requirements
S50 NFV Site Deployment Managing site selection, dynamic VF allocation, and service chaining during NFV deployment in multi-cloud environment Site Selection and VF Allocation Algorithms Development of algorithms for site selection and dynamic VF allocation
S79 Cloud Management Platform Evaluation Deployment Addressing diverging functionality, lack of standard benchmark, and evaluating CMPs during multi-cloud container deployment Standardized Output Formats and Evaluation Criteria Establishment of standardized output formats and evaluation criteria
S37 Standardized Requirement Deployment Standardizing the expression of application requirements during multi-cloud container deployments Application-Level Resource Managers for Multi-Cloud Emergence of application-level resource managers supporting multi-cloud deployment
S42 Heterogeneous Node Deployment Managing the deployment and control of heterogeneous nodes in multi-cloud environment Technology-Independent Multi-Level Adaptation Utilization of MicroCloud and TOSCA library for technology independence and fine-grained multi-level adaptation
S52 Deployment Plan Generation Identifying suitable cloud services and generating deployment plans for multi-cloud environment Cloud Services Selection and Deployment Planning Utilization of the MUSA Decision Support Tool and Deployer tool for cloud services selection and deployment planning
S57 TOSCA-based Deployment Implementing TOSCA-based container deployment across multiple cloud providers Brokered Multi-Cloud Deployment with TOSCA Adoption of a brokered approach for multi-cloud deployment with TOSCA
S58 Standardization Deployment Addressing compatibility issues and lack of standardization during multi-cloud container deployment Proxy Cloud Virtualization for OpenStack Framework mapping other clouds as sub-clouds to OpenStack with Proxy Cloud Virtualization layer
@@
S20 Vendor Lock-in Prevention Deployment Developing strategies to prevent vendor lock-in and ensure modular and loosely-coupled deployment in multi-cloud environment Cloud Modelling Framework Introduction of a Cloud Modelling Framework for specifying provisioning and deployment
S25 Security Deployment in Multi-container Environment Addressing security and privacy concerns during multi-container deployment on the same OS kernel in multi-cloud environment Cross-Container Isolation Enforcing cross-container isolation
S34 Security Deployment Ensuring integrated security during the deployment of multi-cloud applications Security Control Specification and Deployment Framework Utilization of the MUSA framework for specifying, deploying, and assessing security controls
S72 Virtual Network Security Deployment Providing secure virtual networks during multi-cloud container deployment Network Virtualization Platform for Multi-Cloud Adoption of Sirius, a network virtualization platform for multi-cloud environment
S83 Malicious Service Management Deployment Managing malicious services and supporting multi-cloud communities formation during container deployment Collusion-Resilient Trust Aggregation Technique Implementation of a trust aggregation technique resilient to collusion attacks

### v2

@@ 3.6. Automation Challenges and Solution Framework (RQ4 and RQ5)
(Alonso et al., 2019) Automating Multi-Cloud Deployment Monitoring Adopting Multi-Cloud Strategy
(Deng et al., 2022) Optimizing Container Placement Implementing Intelligent Placement
(Zhang et al., 2021) Automating Runtime Application Switching Deploying Service Driver Model
(Pham and Pham, 2015) Enabling Fine-Grained Component Automation Replicating and Migrating Services
(Barna et al., 2017) Streamlining DevOps Automation Developing with DevOps Approach
(Gupta et al., 2021) Automating Service Deployment Using Automated Scripts
(Sheridan et al., 2017) Securing Deployment Time Securing with Automation Tools
(Mittal and Risco-Martín, 2017) Automating Distributed Simulations Adopting DevOps Methodologies
(Al-shammari and Alwan, 2018) Improving Software Deployment Speed Optimizing Container Images
(Lorenzo et al., 2019) Automating Container Startup Profiling Container Execution
(Pandey et al., 2022) Dynamically Scaling Resources Automating Scaling with CSP Capability
(Pahl, 2015) Scaling Container Deployment Scaling with Kubernetes
(Saif et al., 2023) Provisioning Elastic Resources Provisioning Resources Elastically
(Bhamare et al., 2015) Automating VF Placement Optimizing VF Placement
(Guerrero et al., 2018a) Managing Container and VM Deployment Orchestrating Multi-Cloud Deployments
@@
(Ramos et al., 2021) Handling Resource Overbooking Detecting Overbooking with ML
(Cui et al., 2016) Managing Virtual Clusters Managing with ClaaS Model
(Altran et al., 2022) Scheduling Containerized Applications Scheduling with Kubernetes
(Shi et al., 2020) Selecting Optimal VM Types Using GA-based Algorithms
(Baresi et al., 2016) Automating Resource Commissioning Adopting MicroCloud Architecture
(Dreibholz et al., 2019) Supporting Cloud Federation Automation Optimizing with Middleware Platform
(Trihinas et al., 2015) Automating Elastic Resource Detection Automating with JCatascopia
(Buyya and Barreto, 2015) Automating Resource Acquisition Managing Resources with Aneka
(Mfula and Nurminen, 2018) Enforcing SLA Compliance Testing for Compliance
@@
(Almeida et al., 2014) Safeguarding Data Integrity in Migration Maintaining Integrity Autonomously
(Slominski et al., 2015) Migrating Legacy Web Applications Reusing Architectural Patterns
(Kiss, 2018) Porting Applications to Cloud Using CloudSME Platform
Category 5: Testing and Benchmarking Automation
(Verdugo et al., 2017) Automating User-Oriented Testing Virtualizing with Lightweight Containers
(Jha et al., 2019) Streamlining Benchmarking Orchestrating Docker Benchmarking
(Naik, 2016a) Enhancing Automation Dependability Enhancing Dependability with CI
@@
(Gattobigio et al., 2022) Standardizing Automation Processes Integrating Service Mesh
(Lazuka et al., 2022) Optimizing Cloud Provider Selection Selecting Optimal Providers
(Ouyang et al., 2022) Unifying Microservices Deployment Adopting Service Chain Models
(Harwalkar et al., 2019a) Addressing Cloud Interoperability Managing with Multi-Cloud Harmony
(Pellegrini et al., 2017) Standardizing Cloud Applications Standardizing Interfaces
(Pai and Kumar, 2021) Automating Package Creation and Scaling Automating Package Frameworks
@@
(Bolivar et al., 2018) Selecting 5G Ecosystem Components Utilizing Open-Source Components
(Rahman and Lama, 2019) Managing End-to-End Tail Latency Modeling Performance Efficiently
(Serhiienko and Spillner, 2018) Handling Distributed Application Complexity Managing Distributed Applications
@@
(Alaluna et al., 2017) Addressing Network Complexity Managing Network Complexity
(Wahab et al., 2016) Improving Social Network Service Discovery Automating Service Discovery
(Pham et al., 2015) Administering Complex Applications Building Autonomic Systems
Resource Management Automation category reports on challenges and solutions related to automating the management of resources in multi-cloud environment. Out of the 121 studies reviewed, 9 specifically addressed these challenges and provided appropriate solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)). These challenges include resource overbooking management, resource management for containerized applications, and resource commissioning and decommissioning. The proposed solutions include machine learning-based resource overbooking detection for efficiently managing service containers across multiple cloud platforms. To manage containerized applications, custom Kubernetes schedulers, such as label-affinity schedulers, were proposed to optimize orchestration and allocation. Additionally, a GA-based algorithm was suggested to reduce the search space of VM types in data centers, simplifying the resource allocation problem within multi-cloud environment. For resource commissioning and decommissioning, proposals such as the MicroCloud architecture enabled fine-grained resource allocation and coordinated adaptation workflows. The solutions also addressed cloud federation and inter-platform portability challenges, utilizing middleware platforms for security enhancement, cost management, and performance optimization. The JCatascopia automated modular monitoring framework was introduced to monitor runtime configurations and detect elasticity actions. Finally, SLA compliance testing automation was proposed to manage resource allocation processes and enforce corrective actions, ensuring that service level agreements are met across cloud environments.
Data and Application Migration Automation: category reports challenges and solutions related to the automation of testing and benchmarking processes for containerized applications in multi-cloud environment. Among the 121 studies reviewed, 3 identified these challenges along with their proposed solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)). The main challenges include ensuring data integrity during the migration process and enabling the smooth migration of legacy Web applications to cloud services. The proposed solutions involve introducing autonomous management cycles to reduce complexity and ensure effective deployment across multi-cloud environment. Other solutions include reusable architectural patterns to accelerate the modernization of legacy applications in containerized multi-cloud environment. Additionally, the CloudSME Simulation Platform was highlighted for its ability to accelerate the porting of existing applications to cloud infrastructures, with support for more complex tasks such as billing and resource management.
Testing and Benchmarking Automation category highlights the challenges and solutions related to automating the testing and benchmarking processes for containerized applications in multi-cloud environment. Out of the 121 studies reviewed, 3 specifically addressed these challenges and provided relevant solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)). The main challenges revolve around automating user-oriented testing and benchmarking across various cloud platforms. The solutions range from just-in-time deployment and streamlined multicomponent prototyping to rapid testing, enabled by container-based lightweight virtualization. Additionally, Smart Docker Benchmarking Orchestrators were proposed for automating benchmarking in multi-cloud environment, thereby enhancing the overall efficiency of testing procedures. The application of Computational Intelligence (CI) was also suggested to improve dependability in automated testing for containerized multi-cloud systems.
Standardization and Interoperability Challenges category reports on the challenges and solutions related to standardization and interoperability in the automation of containerized applications across multiple cloud environments. Out of the 121 reviewed studies, 6 specifically identified these challenges and proposed corresponding solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)). The main challenges involve standardizing automation processes, selecting cloud providers, and addressing the lack of unified templates for microservices deployment. The proposed solutions include service mesh networks and multi-cloud cluster federation to standardize the automation processes for containerized applications. AI-driven tools for optimal cloud provider selection were also introduced, which help in selecting the most suitable provider and node based on factors such as cost-effectiveness and runtime efficiency in multi-cloud environment. To address the lack of common templates for microservices, a Service Model approach was proposed to automate the deployment of microservices across multiple cloud platforms, thereby reducing operational complexity. For cloud interoperability, an AI-driven platform, such as multi-cloud Harmony, was developed to facilitate seamless data transfer and communication between disparate cloud environments. In terms of cloud application standardization, the solutions emphasized promoting the adoption of global standards and interfaces to improve interoperability between cloud service providers and middleware platforms. Finally, automation frameworks for package creation and scaling, such as those using machine-readable definition files, were proposed to efficiently manage both virtual machines and Docker containers across multi-cloud environment.
Application and Service Management category reports challenges and solutions related to the automation of application and service management processes in containerized multi-cloud environment.Out of the 121 reviewed studies, 2 specifically addressed these challenges and proposed appropriate solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)). The key challenges include the selection of 5G ecosystem components and end-to-end tail latency management in microservice architectures. The proposed solutions involve using open-source software components to customize the orchestration of containers and virtual machines in modern multi-cloud deployments. Additionally, performance modeling approaches were suggested to evaluate the performance of container-level and VM-level data in multi-cloud environment, enabling more efficient latency management and optimized overall performance.
Runtime and Service Discovery category organizes the challenges and solutions related to runtime management and services discovery in multi-cloud environment, particularly in managing the complexity of network control and administration for containerized applications. Out of the 121 reviewed studies, 4 specifically tackled these challenges (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)). The proposed solutions include automating network management to address the increasing complexity of distributed applications across multi-cloud topologies. Additionally, automated discovery and selection algorithms were proposed to tag and manage social network services within containerized multi-cloud systems. Autonomic computing systems were introduced to handle the growing complexity of applications by facilitating knowledge transfer, enabling more effective containerization management through automation.
@@ 3.7. Deployment Challenges and Solution Framework (RQ4 and RQ5)
Table LABEL:tab:Deployment-solutions presents a comprehensive overview of the challenges and corresponding solutions faced during the deployment of containerized applications in multi-cloud environment. As organizations increasingly adopt containerization and multi-cloud strategies to enhance scalability, flexibility, and resource utilization, they encounter various obstacles that require effective solutions. The figure consists of nine categories, each representing a specific deployment challenge, along with the corresponding challenge description and its solution. These categories encompass a wide range of issues, including deployment complexity and Orchestration, access and communication control, security and compliance management, multi-cloud deployment coordination and Integration, monitoring and scalability challenges, microservices architecture and containerization, and network connectivity and hybrid cloud integration challenges.
Deployment Complexity and Orchestration category outlines various challenges and solutions related to the complexity and orchestration of container deployment in multi-cloud environment. Out of the 121 studies reviewed, 18 focused on these specific challenges and their respective solutions. Key challenges include Performance Testing Deployment, where standardized benchmarking and automation ensure consistency in performance testing. Deployment Validation is addressed by the Testing Process Management System, which enhances validation to ensure all criteria are met before deployment. Cost-Effective Deployment is achieved through Containerization and Microservice Architecture, enabling agile and efficient deployment. For Resource Management, a User Preference-Based Resource Brokering system optimizes resource allocation across clouds. Additional prominent solutions include orchestration using TOSCA, automation tools like Jenkins and Terraform, and Blockchain-based SLA Monitoring for immutable and auditable SLA tracking. Dynamic Adaptation is handled through Context-Aware Resource Allocation, while AWS Wavelength for Edge Cloud Integration addresses regional infrastructure consistency issues.
@@
Access and Communication Control category organizes challenges and solutions related to maintaining secure connectivity in multi-cloud environment and managing user access control. Out of the 121 reviewed studies, 8 focused on these challenges. These studies report solutions include ensuring secure access to resources, secure communication between containers, and maintaining secure connectivity across multiple clouds. The solutions involve the development of encryption and authentication mechanisms to protect data and ensure authorized access. Secure communication protocols were recommended for interactions within containers and across cloud environments. It was also suggested that security measures should be provider-agnostic, ensuring that they are not tied to any single cloud provider.
@@
Security and Compliance Management in Multi-cloud Deployment category focuses on security and compliance challenges and their solutions faced during multi-cloud container deployment. One challenge, Vendor Lock-in Prevention Deployment, aims to prevent vendor lock-in and ensure modular and loosely-coupled deployment in multi-cloud environment. The solution proposed is the Cloud Modelling Framework, which introduces a framework for specifying provisioning and deployment to enhance compatibility with existing ACSs and cloud solutions. Another challenge in this category is Security Deployment in Multi-container Environment, which addresses security and privacy concerns during multi-container deployment on the same OS kernel in multi-cloud environment. The solution is Cross-Container Isolation, which enforces cross-container isolation to enhance security.
@@
Table 13. Identified deployment challenge and solution for containerized applications in multi-cloud environment
ID Challenge Solution
@@
(Almeida et al., 2014) Performance Testing Deployment Standardized Benchmarking and Automation
(Verdugo et al., 2017) Deployment Validation Testing Process Management System
(Jha et al., 2019) Cost-effective Deployment Containerization and Microservice Architecture
(Pandey et al., 2022) Resource Management Deployment User Preference-Based Multi-Cloud Resource Brokering
(Bhamare et al., 2015) NFV Site Deployment Site Selection and VF Allocation Algorithms
(Serhiienko and Spillner, 2018) Cloud Management Platform Evaluation Standardized Output Formats and Evaluation Criteria
(Dimitrijevic et al., 2019) Standardized Requirement Deployment Application-Level Resource Managers for Multi-Cloud
(Baresi et al., 2016) Heterogeneous Node Deployment Technology-Independent Multi-Level Adaptation
(Casola et al., 2017) Deployment Plan Generation Cloud Services Selection and Deployment Planning
(Tricomi et al., 2017) TOSCA-based Deployment Brokered Multi-Cloud Deployment with TOSCA
(Sitaram et al., 2018) Standardization Deployment Proxy Cloud Virtualization for OpenStack
(Barletta et al., 2024) Criticality Constraints in Kubernetes Deployment Kubernetes for Critical Deployments
(da Silva et al., 2023) Automation Tools Deployment Issue Accurate Cost and Resource Tracking
(Pandi et al., 2023) Multi-cloud Deployment Challenge Jenkins for Multi-cloud Automation
(Sofia et al., 2024) Decentralized App Deployment Issue CODECO for Edge-Cloud Deployment
(Neeraj et al., 2023) SLA Monitoring Deployment Complexity Blockchain for SLA Data Integrity
(Petrovski and Gusev, 2024) Container and Serverless Architecture Deployment Issue AWS ECS and Lambda Deployment
(Solayman and Qasha, 2023) Heterogeneous Environment Deployment Challenge Edge-cloud Orchestration Management
(Sofia et al., 2023) Dynamic Adaptation Deployment Issue Context-aware Resource Allocation Strategies
(Santos et al., 2023) Regional Infrastructure Deployment Consistency Issue AWS Wavelength for Edge Cloud Integration
(Roy and Kalafatis, 2023) Robotic App Deployment Challenge Hybrid Edge-Cloud Architecture
(Alonso et al., 2023a) Multi-cloud Deployment Consistency Problem Multi-cloud Deployment Efficiency Models
(Harzenetter et al., 2023) Cross-Technology Deployment Integrity Issue Workflow Automation for Cross-Technology Deployment
(Huang and Wang, 2023) Industrial Robot Control System Deployment Problem Seamless Control between Cloud and Robots
(Nagarajan et al., 2023) Efficient Task Deployment across Clouds Hybrid Approach for Resource Allocation
@@
(Ferry et al., 2014) Vendor Lock-in Prevention Deployment Cloud Modelling Framework
(Gao et al., 2017) Security Deployment in Multi-container Environment Cross-Container Isolation
(Rios et al., 2017) Security Deployment Security Control Specification and Deployment Framework
(Alaluna et al., 2017) Virtual Network Security Deployment Network Virtualization Platform for Multi-Cloud
(Wahab et al., 2016) Malicious Service Management Deployment Collusion-Resilient Trust Aggregation Technique
(Barati et al., 2023) Secure Personal Data Deployment Challenge Docker, Kubernetes, Blockchain for GDPR Compliance
(Fugkeaw, 2023) IAM Deployment across Cloud Platforms Blockchain for Decentralized IAM Deployment
(Wang et al., 2023) Non-intrusive Malware Detection Deployment DockerWatch for Non-intrusive Container Integration